1. Determine the absolute pressure at the bottom of a lake that is 30 meters deep.
1
Expert's answer
2013-01-24T09:14:01-0500
The water pressure at the bottom of a lake that is 30 meters deep according to the Pascal's law is

P(water) = pgh = 1000[kg/m³]·9.8[m/s²]·30[m] = 294000[N/s²] = 294000 [Pa].

The usual atmospherical pressure is

P(atm) = 10^5[Pa].

Therefore, the absolute pressure at the bottom of a lake is

P(total) = P(water) + P(atm) = 294000[Pa] + 100000[Pa] = 394000[Pa].